OpenStudy (anonymous):

Zoe A = 10000(1.06)^8. Ned A = 8000 (1+i)^8. They both end up with the same amount at the end.Whats value for i.? So heres what i got ... Zoe A= 10000(1.06)^8 A= 15938.5 Ned. 15938.5=8000(1+i)^8 15938.5/8000=(1+i)^8 1.99^(1/8)= 1+i 1.99^(1/8)-1=i 0.089x100=i 8.9% = i Ned - A = 8000 (1+8.9)^8 Doesnt add up to 15938.5... what am i doing wrong?

OpenStudy (anonymous):

I figured it out myself i was rounding it that was my issue. 23) Sachi. A= 10000(1.06)^8 A= 15938.5 Nari. 15938.5=8000(1+i)^8 15938.5/8000=(1+i)^8 1.99^(1/8)= 1+i 1.99^(1/8)-1=i 0.0899827=i